The birth weight in the United States has a standard deviation of σ = 575. A random sample of n = 30 birth weights give ¯x = 3189. (a) Find a 90% confidence interval and interpret it. [3] (b) Suppose now σ is unknown but a sample size of n = 15 is taken yielding a sample standard deviation of s = 538. Find the 90% confidence interval for µ again. (c) Suppose σ is unknown again but a sample size of n = 255 is taken yielding a sample standard deviation of s = 582. Find the 90% confidence interval for µ again
Answer #1

a)

sample mean 'x̄= 3189
sample size   n= 30
population std deviation σ= 575
std error 'σx=σ/√n= 104.9801569
for 90% CI, value of z= 1.644853627
margin of error E=z*std error    = 172.6769918
lower bound=sample mean-E = 3016.323008
Upper bound=sample mean+E = 3361.676992
from above 90% confidence interval for population mean =(3016.323,3361.6769)

we are 90% confidence that population mean is between in the above interval.

b)

sample mean 'x̄= 3189
sample size   n= 15
sample std deviation s= 538
std error 'sx=s/√n= 138.91
for 90% CI; and 14 df, value of t= 1.761310136
margin of error E=t*std error    = 244.665357
lower bound=sample mean-E = 2944.334643
Upper bound=sample mean+E = 3433.665357
from above 90% confidence interval for population mean =(2944.3346,3433.6653)

c)

sample mean 'x̄= 3189
sample size   n= 255
sample std deviation s= 582
std error 'sx=s/√n= 36.446
for 90% CI; and 254 df, value of t= 1.650874791
margin of error E=t*std error    = 60.16820154
lower bound=sample mean-E = 3128.831798
Upper bound=sample mean+E = 3249.168202
from above 90% confidence interval for population mean =(3128.8317,3249.1682)